# On the Identifying Power of Generalized Monotonicity

### v2

@@ 3.2 Identifying Power of Generalized Monotonicity
For example, $D$ might represent quantity of some treatment, and $Z$ might represent levels of subsidy for the treatment. The restriction in (19) is equivalent to the monotonicity assumption considered in Angrist and Imbens (1995). See Vytlacil (2006) for the connection between this restriction and ordered discrete-choice selection models. Without loss of generality, let $\mathcal{D}=\{0,\ldots,\bar{D}\}$. In this case, (3) is satisfied for $d\in\{0,\bar{D}\}$ for all $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$; to see this, take $z^{*}(0)=\min\{\mathcal{Z}\}$ and $z^{*}(\bar{D})=\max\{\mathcal{Z}\}$. By a straightforward modification of the arguments underlying Theorem 3.2, one can show that the identified sets for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$ are given by (5) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. The ordered monotonicity assumption in (19) therefore has no identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$. In contrast, (3) need not hold for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$ and $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. In Appendix B.3, we show there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We further show $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ for the same $P$. Thus, the ordered monotonicity assumption in (19) does have identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$.
@@ Appendix A Proofs of Main Results
The necessity of (4) has been proved in the main text. On the other hand, fix a $d\in\mathcal{D}$, suppose (4) holds for some $z\in\mathcal{Z}$. Fix a particular $z^{\ast}\in\mathcal{Z}^{*}(d,Q)$ that satisfies (3). If $z=z^{\ast}$ then of course Assumption 2.2 holds for $z$. Suppose $z\neq z^{\ast}$. Then,
@@ A.3.1 Auxillary Results
To present the proof of Theorem 3.2, we first introduce some further notation. Recall from the discussion in Section 2 that if $Q$ satisfies Assumption 2.1, then $P=QT^{-1}$ if and only if
@@
Let $r_{j}^{o}$ denote the $(j+1)$th entry of $r^{o}$ and $r_{j}^{t}$ denote the $(j+1)$th entry of $r^{t}$. In this notation,
$Q\{Y_{d}=y,D_{z}=d\}=\sum_{(r^{o},r^{t}):r^{o}_{d}=y,r^{t}_{z}=d}q(r^{o},r^{t})~{},$
where
@@
Thus, if $Q$ satisfies Assumption 2.1, then $P=QT^{-1}$ if and only if
@@
We note Assumption 2.2 can be expressed as restricting $q(r^{o},r^{t})=0$ for certain $r^{t}$ and $r^{o}$. In particular, if for some $d\in\mathcal{D}$, $r^{t}_{z^{*}(d)}\neq d$ while $r^{t}_{z^{\prime}}=d$ for some $z^{\prime}\neq z^{*}(d)$, then $q(r^{o},r^{t})=0$ for all $r^{o}$. Let $R^{t}$ denote the complement of that set of treatment response types, or equivalently the set of treatment response types that may occur with positive probability under Assumption 2.2:
$R^{t}=\{r^{t}\in\mathcal{D}^{|\mathcal{Z}|}~{}:~{}r^{t}_{z^{*}(d)}\neq d~{}\mbox{implies}~{}r^{t}_{z^{\prime}}\neq d~{}\forall~{}d\in\mathcal{D},z^{\prime}\in\mathcal{Z}\}~{}.$ (21)
If Assumption 2.2 is satisfied, then $Q\{(D_{z}:z\in\mathcal{Z})\in R^{t}\}=1$.
@@
For a given treatment response type $r^{t}$, $\mathcal{N}(r^{t})$ is the set of treatments for which that treatment response type is a never-taker, and $\mathcal{N}(r^{t})^{c}$ is the set of treatments for which that treatment response type will take that treatment for some value of $z$. Using this notation, partition outcome and treatment response types $(r^{o},r^{t})$ as $(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ where
@@
For a given treatment response type $r^{t}$, $r^{o}_{n}(r^{t})$ are those outcomes that are never observed for that response type, and $r^{o}_{c}(r^{t})$ are the remaining outcomes that are observed given some potential value of $Z$. For example, for any $r^{t}\in R^{t}$ with $R^{t}$ defined by (21),
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{*}(d)}\neq d\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{*}(d)}=d\}~{},$
@@
$r^{o}_{n}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{*}(d)}\neq d)~{},$
$r^{o}_{c}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{*}(d)}=d)~{}.$
For notational convenience, further define $q(r^{o}_{c}(r^{t}),r^{t})$ as $q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ summed over $r^{o}_{n}(r^{t})$:
@@
Using this notation, we have the following lemma that asserts whether $q(r^{o},r^{t})$ satisfies (20) depends only on $q(r^{o}_{c}(r^{t}),r^{t})$.
@@
Suppose $q$ satisfies (20). Then, $q^{*}$ satisfies (20) if, for each $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
$q^{*}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})~{}.$
@@ A.3.2 Proof of the Theorem
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. For each $d\in\mathcal{D}$,
$\mathbb{E}_{Q}[Y_{d}]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$
$=\beta_{d|z^{*}(d)}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]~{},$
@@
$\mathbb{E}[Y_{d}]\in[\beta_{d|z^{*}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)}),~{}\beta_{d|z^{*}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{*}(d)})]~{},$
@@
Let $q$ denote latent probabilities corresponding to a fixed $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$, which is nonempty by assumption. We now show that for any value in (5) there exists $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$ for which $(\mathbb{E}_{Q^{\ast}}[Y_{j}]:j\in\mathcal{D})$ equals the prescribed value. To do so, we show that, for each $\theta$ in the right-hand side of (5), we can construct $q^{\ast}$ that 1) is nonnegative and has $\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})$, 2) satisfies (20) and hence $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$, and 3) satisfies $\theta({Q^{\ast}})=\theta$.
@@
Fix some $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ to be specified below, we now construct an alternative $q^{\ast}$ as follows. For each treatment response type $r^{t}$, let $q^{\ast}_{0}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}$. Let $K(r^{t})=|\mathcal{N}(r^{t})|$. Note that if $K(r^{t})=0$, then $\mathcal{N}(r^{t})=\emptyset$, so $r^{t}_{z^{*}(d)}=d$ for all $d\in\mathcal{D}$ under $r^{t}$. For such an $r^{t}$, we set $q^{\ast}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$. If $K(r^{t})\geq 1$, enumerate $\mathcal{N}(r^{t})$ as $\{j[1],...,j[K(r^{t})]\}$, and for $k=1$ to $K(r^{t})$, define $q_{k}^{\ast}$ iteratively as follows:
$\begin{split}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})&=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{*}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})&=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}\\ q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})&=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{*}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})~{},\end{split}$ (22)
@@
for all $r^{o}_{-j[k]}$, where we partition $r^{o}=(r^{o}_{-j[k]},r^{o}_{j[k]})$. With this construction, the marginal distribution of $Y_{j[k]}$ for treatment response type $r^{t}$ is only modified in step $k$. In particular, for each fixed $k$, for $\ell\leq k-1$, for $y\in\mathcal{Y}$,
$\sum_{r_{-j[k]}^{o}}q^{*}_{\ell}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q^{*}_{0}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})~{},$ (23)
which together with (22) implies
$\mathbb{E}_{Q^{\ast}_{k}}[Y_{j[k]}\mathbbm{1}\{(D_{z}:z\in\mathcal{Z})=r^{t}\}]$ $=\sum_{y\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}yq^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})$
$=(\alpha_{j[k]}y^{U}+(1-\alpha_{j[k]})y^{L})\sum_{r^{o}}q(r^{o},r^{t})~{}.$

### v3

@@ 3.2 Identifying Power of Generalized Monotonicity
For example, $D$ might represent quantity of some treatment, and $Z$ might represent levels of subsidy for the treatment. The restriction in (13) is equivalent to the monotonicity assumption considered in Angrist and Imbens (1995). See Vytlacil (2006) for the connection between this restriction and ordered discrete-choice selection models. Without loss of generality, let $\mathcal{D}=\{0,\ldots,\bar{D}\}$. In this case, (3) is satisfied for $d\in\{0,\bar{D}\}$ for all $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$; to see this, take $z^{\ast}(0)=\min\{\mathcal{Z}\}$ and $z^{\ast}(\bar{D})=\max\{\mathcal{Z}\}$. By a straightforward modification of the arguments underlying Theorem 3.2, one can show that the identified sets for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$ are given by (5) for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. The ordered monotonicity assumption in (13) therefore has no identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{0}]$ and $\mathbb{E}_{Q}[Y_{\bar{D}}]$. In contrast, (3) need not hold for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$ and $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. In Appendix B.3, we show there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We further show $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ for the same $P$. Thus, the ordered monotonicity assumption in (13) does have identifying power beyond instrument exogeneity for $\mathbb{E}_{Q}[Y_{d}]$ for $d\in\mathcal{D}\setminus\{0,\bar{D}\}$.
@@ Appendix A Proofs of Main Results
The necessity of (4) has been proved in the main text. On the other hand, fix a $d\in\mathcal{D}$, suppose (4) holds for some $z\in\mathcal{Z}$. Fix a particular $z^{\ast}\in\mathcal{Z}^{\ast}(d,Q)$ that satisfies (3). If $z=z^{\ast}$ then of course Assumption 2.2 holds for $z$. Suppose $z\neq z^{\ast}$. Then,
@@
This section is organized as follows. In Section A.3.1, we introduce additional notation that is helpful in formally proving our result, including defining subgroups of individuals, called treatment response types, who are defined by what treatment they would take at each value of the instrument. Because all variables are discrete, we will directly work with the probability mass function. We derive a lemma that characterizes a sufficient condition for a given distribution of potential outcomes and treatments $Q$ to rationalize the distribution of the data $P$. The lemma states that whether a $Q$ that satisfies Assumption 2.1 rationalizes $P$ depends on the probability of each treatment response type and on the probability of that type’s potential outcomes corresponding to treatments they would choose for some value of the instrument (so that they “comply with” this treatment at least for some values of the instrument), but does not depend on the probability of that type’s potential outcomes corresponding to treatments they would not choose for any value of the instrument (so that they are “never-takers” of this treatment). If $Q$ satisfies Assumptions 2.1 and 2.2, then the set of treatments for which a treatment response type is a “never-taker” are precisely the set of treatments that they would not take even when maximally encouraged to do so. Therefore, the implication of the lemma is that if $Q$ satisfies Assumptions 2.1 and 2.2 and rationalizes $P$, then any other $Q^{\ast}$ satisfying Assumption 2.1 and 2.2 will also rationalize $P$ if $Q$ and $Q^{\ast}$ differ only in the probability of potential outcomes corresponding to treatments that a given response type would not take even when maximally encouraged to do so.
In Section A.3.2, we use the notation and lemma introduced in Section A.3.1 to prove the theorem. Let $\mathbf{Q}$ satisfy the assumptions of the theorem. We first show that $\Theta_{0}(P,\mathbf{Q})$ is a subset of the bounds in (5). We then show that the bounds in (5) are a subset of $\Theta_{0}(P,\mathbf{Q})$ using the following proof strategy. By assumption, $\Theta_{0}(P,\mathbf{Q})$ is non-empty, so that there exists a distribution $Q\in\mathbf{Q}$ that rationalizes $P$. For each value $\theta_{0}$ in the bounds of (5), we construct an alternative distribution $Q^{\ast}\in\mathbf{Q}$ such that $\theta(Q^{\ast})=\theta_{0}$ with $Q$ and $Q^{\ast}$ differing only in the probability of outcomes corresponding to treatments that a given response type would not take even when maximally encouraged to do so. That the constructed $Q^{\ast}$ lies in $\mathbf{Q}$ follows from the assumption that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 and that $Q$ and $Q^{\ast}$ have the same distribution of potential treatment choices with $Q\in\mathbf{Q}$. That the constructed $Q^{\ast}$ rationalizes $P$ follows from $Q\in\mathbf{Q}$ and the previously described lemma. That we are able to construct such a $Q^{\ast}$ with $\theta(Q^{\ast})=\theta_{0}$ for every $\theta_{0}$ in the bounds of (5) establishes that the bounds (5) are a subset of $\Theta_{0}(P,\mathbf{Q})$, completing the proof.
@@ A.3.1 Auxillary Results
To present the proof of Theorem 3.2, we first introduce some further notation. Because all variables are discrete, we will directly work with the probability mass function. Recall from the discussion in Section 2 that if $Q$ satisfies Assumption 2.1, then $P=QT^{-1}$ if and only if
@@
Because all variables are discrete, we define the probability mass function as
@@
For the rest of the proof, without loss of generality, we suppose $\mathcal{D}=\{0,1,\dots,|\mathcal{D}|-1\}$ and $\mathcal{Z}=\{0,1,\dots,|\mathcal{Z}|-1\}$. Let $r_{j}^{o}$ denote the $(j+1)$th entry of $r^{o}$ and $r_{j}^{t}$ denote the $(j+1)$th entry of $r^{t}$. In other words, $r_{j}^{o}$ denotes the value of the potential outcome for the outcome response type $r^{o}$ when the treatment equals $j$, and $r_{j}^{t}$ denotes the value of the potential treatment for the treatment response type $r^{t}$ when the instrument equals $j$. In this notation, if $Q$ satisfies Assumption 2.1, then it follows from (2) that $P=QT^{-1}$ if and only if
@@
For a given treatment response type $r^{t}$, $\mathcal{N}(r^{t})$ is the set of treatments for which that treatment response type is a “never-taker,” and $\mathcal{N}(r^{t})^{c}$ is the set of treatments for which that treatment response type will “comply with” the treatment for some value of $z$. Using this notation, partition outcome and treatment response types $(r^{o},r^{t})$ as $(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ where
@@
For a given treatment response type $r^{t}$, $r^{o}_{n}(r^{t})$ are those outcomes that are never observed for that response type, and $r^{o}_{c}(r^{t})$ are the remaining outcomes that are observed given some value of $Z$. Here, the subscripts $n$ and $c$ stand for “never-taker” and “complier.”
Remark A.1.
Here we illustrate how our notation specializes under Assumption 2.2. Note Assumption 2.2 can be expressed as restricting $q(r^{o},r^{t})=0$ unless the treatment response type $r^{t}$ satisfies the condition therein; in other words, it restricts the support of the treatment response type. In particular, if for some $d\in\mathcal{D}$, $r^{t}_{z^{\ast}(d)}\neq d$ while $r^{t}_{z^{\prime}}=d$ for some $z^{\prime}\neq z^{\ast}(d)$, then $q(r^{o},r^{t})=0$ for all $r^{o}$. For any $r^{t}$ in the support,
$\mathcal{N}(r^{t})$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{\ast}(d)}\neq d\}~{},$
$\mathcal{N}(r^{t})^{c}$ =$=\{d\in\mathcal{D}~{}:~{}r^{t}_{z^{\ast}(d)}=d\}~{},$
@@
$r^{o}_{n}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{\ast}(d)}\neq d)~{},$
$r^{o}_{c}(r^{t})$ $=(r^{o}_{d}:d\in\mathcal{D},~{}r^{t}_{z^{\ast}(d)}=d)~{}.$
Indeed, $z^{\ast}(d)$ is the instrument that maximally encourages to treatment $d$, so if $r^{t}_{z^{\ast}(d)}\neq d$, then $r^{t}_{z}\neq d$ for all $z\in\mathcal{Z}$. As a result, someone with that treatment response type $r^{t}$ never takes $d$, and hence $d\in\mathcal{N}(r^{t})$. Otherwise, $d\in\mathcal{N}(r^{t})^{c}$, or this person is a “complier” for treatment $d$ at least when $z=z^{\ast}(d)$. The outcome response type $r^{o}$ is then partitioned into $r^{o}_{n}(r^{t})$ and $r^{o}_{c}(r^{t})$ according to whether $d\in\mathcal{N}(r^{t})$ or not.
For notational convenience, we further define the probability mass $q(r^{o}_{c}(r^{t}),r^{t})$ as $q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$ summed over $r^{o}_{n}(r^{t})$:
@@
In defining the probability mass $q(r^{o}_{c}(r^{t}),r^{t})$, we sum over all possible values of $r^{o}_{n}(r^{t})$, because these are the outcomes of treatments that are never taken by the treatment response type $r^{t}$, and hence will not be relevant for the observed data. Using this notation, we have the following lemma that asserts whether $q(r^{o},r^{t})$ satisfies (14) depends only on $q(r^{o}_{c}(r^{t}),r^{t})$. This lemma implies that whether a distribution of potential outcomes and treatments $Q$ that satisfies Assumption 2.1 rationalizes the distribution of the data $P$ depends only on the probability of each treatment response type and the probability of that type’s potential outcomes that would be observed for some value of the instrument.
@@
Suppose $q$ satisfies (14). Then, $q^{\ast}$ satisfies (14) if, for each $r^{t}\in\mathcal{D}^{|\mathcal{Z}|}$,
$q^{\ast}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})~{}.$ (15)
@@ A.3.2 Proof of the Theorem
[…]
We first show that (5) provides valid bounds on $\theta(Q)$ under the stated assumptions, that is, $\Theta_{0}(P,\mathbf{Q})$ is a subset of the bounds of (5). Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. For each $d\in\mathcal{D}$,
$\mathbb{E}_{Q}[Y_{d}]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}\neq d\}]$
$=\beta_{d|z^{\ast}(d)}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}\neq d\}]~{},$
@@
$\mathbb{E}[Y_{d}]\in[\beta_{d|z^{\ast}(d)}+y^{L}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)}),~{}\beta_{d|z^{\ast}(d)}+y^{U}(1-\sum_{y\in\mathcal{Y}}p_{yd|z^{\ast}(d)})]~{},$
@@
We now show that the bounds of (5) are the identified set for $\theta(Q)$, that is, the bounds of (5) are a subset of $\Theta_{0}(P,\mathbf{Q})$. Let $q$ denote latent probabilities corresponding to a fixed $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})$. There exists such a $q$ by the assumption that $\mathbf{Q}_{0}(P,\mathbf{Q})$ is non-empty. We show that for each $\theta_{0}$ in the right-hand side of (5), we can construct an alternative distribution of potential outcomes and treatments $Q^{\ast}$ that is contained in $\mathbf{Q}_{0}(P,\mathbf{Q})$ and for which $\theta(Q^{\ast})=(\mathbb{E}_{Q^{\ast}}[Y_{j}]:j\in\mathcal{D})$ is equal to $\theta_{0}$. In particular, for each $\theta_{0}$ in the right-hand side of (5) we will construct $q^{\ast}$ corresponding to $Q^{\ast}$ that
  1. (a)
    satisfies $\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})$ and hence $Q^{\ast}\in\mathbf{Q}$ because the distribution of $(D_{z}:z\in\mathcal{Z})$ is unchanged and by assumption that $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1,
  2. (b)
    satisfies (15) and hence $P=Q^{\ast}T^{-1}$ due to Lemma A.1, and
  3. (c)
    satisfies $\theta({Q^{\ast}})=\theta_{0}$.
Properties (a) and (b) allow us to conclude that $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$, that is, the constructed distribution is consistent with $P$ and the model $\mathbf{Q}$. These properties will follow from our iterative construction of $q^{\ast}$, which preserves the marginal distribution of potential treatments but modifies the marginal distributions of potential outcomes for outcomes that are never observed for a given treatment response type, that is, correspond to a never-taken treatment for a given treatment response type. Because the marginal distribution of potential treatments is preserved, property (a) follows. Because only the marginal distributions of potential outcomes for never-taken treatments are modified, property (b) follows. Property (c) follows from being able to flexibly modify the marginal distributions of potential outcomes for never-taken treatments, so that any $\theta$ can be achieved.
Part 1: construct $Q^{\ast}$
@@
We now construct an alternative $q^{\ast}$ as follows. Fix some vector of weights $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ to be specified below. For each treatment response type $r^{t}$, let $q^{\ast}_{0}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}$. Let $K(r^{t})=|\mathcal{N}(r^{t})|$ be the number of treatments for which treatment response type $r^{t}$ is a never-taker. Note that if $K(r^{t})=0$, then $\mathcal{N}(r^{t})=\emptyset$, so $r^{t}_{z^{\ast}(d)}=d$ for all $d\in\mathcal{D}$ under $r^{t}$. For such an $r^{t}$, we set $q^{\ast}(r^{o},r^{t})=q^{\ast}_{0}(r^{o},r^{t})=q(r^{o},r^{t})$ for all $r^{o}\in\mathcal{Y}^{|\mathcal{D}|}$.
If $K(r^{t})\geq 1$, enumerate the set of never-taken treatments $\mathcal{N}(r^{t})$ as $\{j[1],...,j[K(r^{t})]\}$, and for $k=1$ to $K(r^{t})$, define $q_{k}^{\ast}$ iteratively as follows:
$\begin{split}q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})&=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{\ast}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})&=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}\\ q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})&=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}q^{\ast}_{k-1}((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})~{},\end{split}$ (16)
for all $r^{o}_{-j[k]}$, where we partition $r^{o}=(r^{o}_{-j[k]},r^{o}_{j[k]})$. Intuitively, in step $k$, for each $r^{o}_{-j[k]}$ and $r^{t}$ we reassign the probabilities of all outcome responses to never-taken treatment $j[k]$ to outcome responses $y^{L}$ and $y^{U}$, splitting between $y^{L}$ and $y^{U}$ according to weight $\alpha_{j[k]}$.
@@
With this construction, the marginal distribution of $Y_{j[k]}$ for treatment response type $r^{t}$ is only modified in step $k$. This statement implies that for each fixed $k$, for step $\ell\leq k-1$, and for any outcome $y\in\mathcal{Y}$,
$\sum_{r_{-j[k]}^{o}}q^{\ast}_{\ell}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q^{\ast}_{0}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})=\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})~{}.$ (17)